Problem 12-21 Firm Valuation

Happy Times, Inc., wants to expand its party stores into the Southeast. In order to establish an immediate presence in the area, the company is considering the purchase of the privately held Joe’s Party Supply. Happy Times currently has debt outstanding with a market value of $200 million and a YTM of 6 percent. The company’s market capitalization is $440 million, and the required return on equity is 11 percent. Joe’s currently has debt outstanding with a market value of $33.5 million. The EBIT for Joe’s next year is projected to be $13 million. EBIT is expected to grow at 8 percent per year for the next five years before slowing to 4 percent in perpetuity. Net working capital, capital spending, and depreciation as a percentage of EBIT are expected to be 7 percent, 13 percent, and 6 percent, respectively. Joe’s has 2.15 million shares outstanding, and the tax rate for both companies is 30 percent.

a. What is the maximum share price that Happy Times should be willing to pay for Joe’s? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Maximum share price           $

After examining your analysis, the CFO of Happy Times is uncomfortable using the perpetual growth rate in cash flows. Instead, she feels that the terminal value should be estimated using the EV/EBITDA multiple. The appropriate EV/EBITDA multiple is 8.

b. What is your new estimate of the maximum share price for the purchase? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Maximum share price           $
Answer #1

A) EBIT at end of 5 years =13(1+0.08)^5 =19.10.

Cash flow at Year 5= 19.10 +.06*19.10(depr)-0.13*19.10(capitalspends)-0.07*19.10(Nwc expense)=0.86*19.10=16,42

Terminal Value = FCF5 (1-T)(1+G)/(r-g)

SO terminal value = (16.42*0.7(1+0.04)/(0.11-0.04) = 170.77

Ans A : 170.77Million for Joe as per this method

Ans B

Lets assume Max value of share price of Joe that Happy should pay =X

EV of Joe = 2.15M*X +Debt-Cash. EBIT of Joe = 13. D&A of Joe =6% of EBIT so EBITDA = 13*1.06=13.78

Cash =13(1+(0.06-0,07-0.13)*0.7=7.64

EV/EBITDA = (2.15X+33.5-7,64)/13.78 = 8 2.15X+25.86=13.78*8=110.24 Solving for X= (110.24-25.86)/2.15 =39.24 per share Happy should pay